OpenStudy (anonymous):

Please help! will give medal. Simplify 9^-2.5 a. 1/243 b. -22.5 c. -3.6 d. -45/18

OpenStudy (anonymous):

\[9^{-2.5}\] Where, we make it into a ratio and make the (-) a positive \[1/9^{2.5}\]

OpenStudy (zzr0ck3r):

\(9^{-2.5}=9^{-2}9^{-\frac{1}{2}}=\frac{1}{9^2}*\frac{1}{\sqrt{9}}=\frac{1}{81*3}=?\)

OpenStudy (anonymous):

9^-2.5=1/243

OpenStudy (anonymous):

1/243?

OpenStudy (anonymous):

Thanks, guys. :)

OpenStudy (anonymous):

@zzr0ck3r Thanks for showing that. That is a much more efficient/correct way of doing it than I was about to do haha